a rectangular garden has the width of 5 feet and a length of 10 feet. if an equal amount is added to both the width and length, the area is increased to 126 square feet. what is this amount?
Answer by addingup(3677)

(5+x)*(10+x) = 126
rewrite:
(x+5)(x+10) = 126
x^2+15x+50 = 126
x^2+15x-76 = 0
(x-4)(x+19) = 0
split in two:
x-4 = 0 or x+19 = 0
x = 4 or x = -19
We are not looking for a negative number, so let's set the -19 aside and try the 4:
(5+4)(10+4) = 126
9*14 = 126 Correct
:
John